OpenStudy (anonymous):

use the co function identities to evaluate expression cos^2 63+cos^2 49+cos^2 27+ cos^2 41

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

pls help

OpenStudy (anonymous):

since \(90-27=63\) you have \[\cos(27)=-\sin(63)\]

OpenStudy (anonymous):

and so \[\cos^2(63)+\sin^2(63)=1\]

OpenStudy (anonymous):

bet the other part gives \(1\) as well by the same reason

OpenStudy (anonymous):

thank you! is there anyway you can explain further. like how did you know which number to use?